23 How long would it take for the price level to double if inflation persisted at...

23

How long would it take for the price level to double if inflation persisted at the following percentages?

Instructions: Enter your answers as whole numbers.

a. 17.5 percent per year:

b. 35 percent per year:

c. 3.5 percent per year:
Answer #1

ANSWER:

A) 17.5%

time taken for price level to double = 70 / inflation per year = 70 / 17.5 = 4 years

B) 35%

time taken for price level to double = 70 / inflation per year = 70 / 35 = 2 years

C) 3.5%

time taken for price level to double = 70 / inflation per year = 70 / 3.5 = 20 years